OpenStudy (anonymous):

pi/4 in degrees? my choices are 30° 135° 225° 45°

OpenStudy (anonymous):

The equation I got pi/4 with is ;\[\cos^{-1} (\frac{ \sqrt{2} }{ 2 }) \]

OpenStudy (anonymous):

@amistre64 @ganeshie8 @ParthKohli @jamielovesmickey @amistre64 @Danielp @brooklyn95 @Hotchellerae21 @Homeworked

OpenStudy (anonymous):

I just need to know how to convert into degrees, or whichever option it should be.

OpenStudy (tkhunny):

Always \(\pi = 180º\)

OpenStudy (anonymous):

it D i just checked

OpenStudy (anonymous):

*its

OpenStudy (anonymous):

Aw yay! yeah because if pi=180° then you divide by 4. That makes sense! thank you so much!

OpenStudy (tkhunny):

Super!! Good work.

OpenStudy (anonymous):

ur welcome:-D

OpenStudy (anonymous):

I wish I could give you both best response. So ill just fan ya both. lol

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

@mrs.purpleswag can you help me with a few more???

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

Ill Dm you!

OpenStudy (anonymous):

????